# ReefNet: A Large-scale, Taxonomically Enriched Dataset and Benchmark for Hard Coral Classification

### v1

@@ 1 Introduction
Coral reefs are among the most biodiverse ecosystems on Earth, providing immense ecological and economic value (Barbier et al., 2011; Spalding et al., 2017). Yet these vibrant habitats are increasingly threatened by anthropogenic pressures, including climate change, overfishing, pollution, and ocean acidification (Cooley et al., 2023; Bellwood et al., 2004; Hughes et al., 2017). As these stressors intensify, coral cover and overall reef resilience decline, underscoring the need for effective conservation and restoration (Brandl et al., 2019; Bellwood et al., 2004). A critical component of reef conservation is habitat mapping and coverage analysis, which typically relies on expert taxonomists to annotate in situ underwater images (Hill & Wilkinson, 2004; Beijbom et al., 2015). However, this dependence on manual labeling severely limits the scale and speed of monitoring (Brandl et al., 2019; Bellwood et al., 2004). Machine learning (ML) offers a path to automation, but progress is impeded by two factors: (i) the scarcity of well-established, ML-ready datasets and benchmarks, and (ii) persistent domain shift, whereby classifiers degrade on sites with different photographic conditions, local assemblages, or label conventions (Belcher et al., 2023; Chen et al., 2021).
CoralNet (Beijbom et al., 2015; Chen et al., 2021), one of the leading annotation platforms, fine-tunes a global model on user-defined “sources.” However, these source-specific classifiers frequently show poor performance on unseen sites owing to variations in taxonomy, imaging protocols, and local reef composition (Williams et al., 2019). ReefCloud.ai (AIMS, 2024) has improved label standardization but remains inaccessible to public researchers. To address these data and accessibility gaps, we introduce ReefNet as a complementary resource, providing a large-scale, publicly available coral reef imagery dataset with point-label annotations taxonomically aligned to the World Register of Marine Species (WoRMS(Board, 2024). Figure 1 shows an overview of the dataset’s structure. ReefNet unifies data from 76 carefully selected CoralNet sources and includes an additional $1.3\text{K}$ images that we collected and annotated from the Al-Wajh lagoon in the Red Sea, comprising 4,6094{,}609 expert annotations of Red Sea hard coral genera, offering a valuable contribution from an understudied biogeographic region. In total, the dataset contains $\sim 925\text{K}$ genus-level scleractinian coral annotations (Table 1). Rigorous manual verification by marine biologists was conducted on 8,962 annotations spanning various sources and labels. This process yielded a high-confidence repository of hard coral annotations across diverse biogeographic regions. In addition to point annotations, we enrich ReefNet with textual descriptions for each genus generated from scanned books (Veron, 2000a; b; Wallace, 1999), enabling language-grounded classification and supporting the development of vision-language models (VLMs).
@@ 2 Related Work
Several large-scale machine-learning datasets have recently emerged to facilitate biodiversity classification. One example is TreeOfLife-10M (Stevens et al., 2023c), which consolidates imagery from iNat2021 (Horn et al., 2021), BIOSCAN-1M (Gharaee et al., 2024), and the Encyclopedia of Life (eol.org), surpassing 10 million images of terrestrial and marine organisms. Despite its scope, underwater taxa remain underrepresented in such general biodiversity repositories, largely because of the substantial logistical challenges associated with acquiring and annotating in situ marine imagery. To address this gap, BenthicNet (Lowe et al., 2024) was introduced, aggregating seafloor images from multiple surveys, including the Seaview Survey Photo-quadrat (González-Rivero et al., 2019) and Reef Life Survey (Edgar et al., 2020), among others (Friedman, 2020). BenthicNet constitutes a major step toward comprehensive marine coverage, but it primarily supports broad-scale benthic habitat labeling using the CATAMI classification scheme (Althaus et al., 2015), and it also supports the WORMS (Board, 2024)taxonomy as it has 887533 annotations of which 287K are hard coral (’Scleractinia’) annotations.(Table 1). Other coral-specific datasets have been proposed to support tasks such as dense segmentation. For instance, CoralVOS and CoralSCOP (Ziqiang et al., 2023; Zheng et al., 2024) offer pixel-level masks but generally categorize corals only at the order level (Scleractinia). Consequently, their utility for fine-grained ecological studies is limited. In contrast, MosaicsUCSD and Eilat (Edwards et al., 2017; Beijbom et al., 2016; Alonso et al., 2019) provide annotations at genus or species levels but are geographically constrained to a single reef site, reducing their ability to generalize across regions. Meanwhile, CoralNet (Beijbom et al., 2015) hosts a large volume of benthic imagery with both human- and machine-generated labels, but does not typically map them to an accurate, globally recognized taxonomy such as the WoRMS. Coralscapes (Sauder et al., 2025) extends benthic classification to a sizeable Red Sea dataset with over 170K polygon annotations, focusing on habitat cover types rather than genus- or species-level coral identification. ReefNet builds on prior efforts while addressing several key limitations, providing one of the most comprehensive collections of expert-verified coral annotations available. These annotations are mapped to the genus level across a diverse, global set of reef sites, addressing challenges like limited taxonomic resolution, geographic bias, and inconsistent labeling standards. By aligning its taxonomy with the World Register of Marine Species (WoRMS), ReefNet ensures consistency and compatibility with ongoing coral biodiversity research, supporting a broad spectrum of ecological and machine learning tasks from habitat classification to genus-level analysis. We provide further clarification of our advantages over CoralNet in appendix A.1.
@@
Table 1: Comparison With Existing Coral Classification Datasets. *Image counts reflect availability as of May 13, 2025. Hard coral/genus-level annotations are not specified for CoralNet, since its labels are not standardized. MosaicsUCSD includes 16 annotated orthomosaics, each from $\sim$1.5K images. CoralSCOP (Zheng et al., 2024) contains 330,144 binary coral/non-coral masks (not limited to hard corals). CoralVOS (Ziqiang et al., 2023) likewise provides only binary coral masks, so a hard coral count is N/A.
@@
CoralNet (Beijbom et al., 2015) World Sparse points Species No 4,524,792* N/A No
CoralSCOP (Zheng et al., 2024) World Masks Order No 41,297 330,144 No
CoralVOS (Ziqiang et al., 2023) 17 sites (South China Sea) Masks Order No 60,456 N/A No
MosaicsUCSD (Edwards et al., 2017) Palmyra Masks Species No 16 44,008 Yes
Eilat (Raphael et al., 2020) Eilat Sparse points Genus No 212 $\sim$12,000 Yes
BenthicNet (Lowe et al., 2024) World
@@
Coralscapes (Sauder et al., 2025) Red Sea (5 countries) Masks N/A No 2,075 N/A No
@@ 3.2 Label Standardization and Mapping
To ensure taxonomic traceability and biological consistency, we mapped annotation labels to the World Register of Marine Species (WoRMS Board (2024)). This step was applied, where applicable, to annotations from the 76 curated CoralNet sources and the Al-Wajh lagoon dataset. Hard coral labels were manually aligned with canonical scientific names and corresponding AphiaIDs [^1] An AphiaID is a unique identifier from the WoRMS database, linking each label to its corresponding entry in the hierarchical taxonomy. As coral taxonomy evolves, AphiaIDs allow users to verify the current validity of labels.. This mapping enabled consistent aggregation of biological entities across sources and ensured compatibility with other biodiversity databases. The initial mapping revealed $\sim$920k annotations at multiple taxonomic levels, consolidated to genus-level hard corals. The dataset covers 44 reef-building hard coral (Order: Scleractinia) genera spanning 20 families. However, this initial label space exhibited inconsistencies in naming, taxonomic granularity, and semantic intent, necessitating further filtering and restructuring.
@@
Label Filtering and Semantic Consolidation. To construct a benchmark dataset suitable for training generalizable AI models, we applied a biologically informed filtering strategy targeting syntactic variation (e.g., “Staghorn coral” vs. Acropora cervicornis) and taxonomic inconsistency (e.g., mixing species-, genus-, and family-level labels). This step was motivated by the inherent difficulty of fine-grained coral identification from imagery alone (Chen et al., 2021; Lowe et al., 2024), which often results in community datasets with inconsistent or imprecise taxonomic labeling.
@@
Figure 3: Geographic Distribution of the Annotations. Sources located within two degrees of Latitude and Longitude are grouped into a single point. One source did not contain any location data and is displayed in Antarctica. Marine Ecoregions of the World are shown in light blue (Spalding et al., 2007).
@@ 3.3.1 Annotation Distribution
The ReefNet dataset comprises 334,162 images, of which 181,223 contain hard coral annotations, totaling 924,626 point annotations. These annotations span 26 marine ecoregions across the tropical belt sensu (Spalding et al., 2007) (see appendix A.2.5 for details). The distribution is uneven: the Hawaiian Ecoregion alone contributes over 221K annotations, followed by the Samoa Islands (202K) and the Mariana Islands (108K). In contrast, regions such as the Southwestern Caribbean, Floridian, and Eastern Brazil are sparsely represented, each with fewer than 8K samples. This geographic imbalance reflects broader disparities in coral monitoring efforts and may influence model generalization to underrepresented areas. Among all annotations, 921,351 are assigned to one of 44 classes, capturing a wide range of coral diversity. The most frequent genera, Porites, Montipora, Acropora, and Pocillopora, dominate shallow reefs globally (Figure 2, Figure 4), while several genera appear in less than 1K annotations, illustrating the dataset’s imbalanced distribution.
@@ 4 ReefNet Benchmarks
A key practical challenge for coral-reef AI is the domain gap encountered when models trained on imagery from one dataset (a “source”) are applied to imagery collected by another dataset. Models trained on datasets from specific locations frequently exhibit poor generalization to ecologically or geographically distinct areas (Williams et al., 2019; Belcher et al., 2023; Wyatt et al., 2025). Variations in camera setups, water clarity, depth, and annotation protocols, even for the same taxa, can degrade performance. To systematically evaluate model robustness and practical applicability in genus-level coral classification, we define two primary benchmarking scenarios: (i) Within-source fine-tuning: Images from a single source (Source A) are partitioned into distinct training, validation, and test subsets. Models are trained on a labeled subset and evaluated on the remainder. Although used by platforms such as CoralNet and ReefCloud, this method requires site-specific manual annotations, limiting scalability. (ii) Cross-source deployment: Models trained on all available images from Source A are directly evaluated on a different source (Source B), without further fine-tuning. This setting assesses generalization across domains, highlighting the model’s ability to overcome domain shifts. ReefNet provides benchmarks for both settings with different training and testing configurations. A quantitative overview of all the splits is in Table 2.
@@ 4.1 ReefNet Within‑Source Split
For the within-source setting, we consider two splits that share an identical label set but with different expert agreement quality. In each split, each source is divided on the image level to prevent information leakage using Multilabel Stratified Shuffle Split (Sechidis et al., 2011; Brady, 2017). The two data‑quality variants considered: (i) Train-S1 / Test-S1, with no source–genus filtration, of 81 % expert agreement, 802,956 training annotations from 70 sources; 9,972 validation annotations from 64 sources; 40,881 test annotations from 69 sources. (ii) Train-S2 / Test-S2, with source–genus filtration, of 92 % expert agreement percentage, 445,985 training annotations from 68 sources; 9,999 validation annotations from 63 sources; 40,881 test annotations from 66 sources.
@@ 5.1 Experimental setup
Evaluated Models. We evaluate three types of models across all benchmarks: (i) Large Scale Pre-trained models fine-tuned on our dataset (Tan & Le, 2019; Chen et al., 2021; He et al., 2016; Liu et al., 2022; Dosovitskiy et al., 2021; Bao et al., 2022; Liu et al., 2021; Touvron et al., 2021), (ii) Vision-Language Models (VLMs) evaluated in a zero-shot setting (Ilharco et al., 2021; Radford et al., 2021; Stevens et al., 2023a; Zhai et al., 2023), and (iii) Multimodal Large Language Models (MLLMs) also evaluated in a zero-shot setting (Bai et al., 2025). Model performance is evaluated using Macro Recall, sometimes referred to as the balanced accuracy score (Pedregosa et al., 2011). Macro Recall is particularly suited for our task as the ReefNet data shows a high degree of class imbalance 2. Additionally, for top-performing models, per-class recall, precision, and F1 scores are provided in appendices( A.4.3A.4.5). Additionally, we fine-tune BioCLIP on our hierarchical dataset, down to the Genus level (referred to as BioCLIP-FT), to investigate its performance on hard coral hierarchical classification, as it reported good performance on other datasets spanning plants, animals, and fungi on TreeOfLife-10M (Stevens et al., 2023c). The details about the training setup for all models have been included in the supplementary material.
@@ 5.2 Finetuning Results
EfficientNet (Tan & Le, 2019) 78.36 82.41 48.01 41.78 69.29 64.59
ResNet (He et al., 2016) 63.91 64.49 37.96 28.49 64.14 51.22
ConvNext (Liu et al., 2022) 77.52 82.88 42.21 47.06 69.19 68.08
BEiT (Bao et al., 2022) 72.18 73.63 45.56 32.17 64.24 62.31
DeiT (Touvron et al., 2021) 74.36 77.40 45.46 40.94 63.19 66.67
Swin (Liu et al., 2021) 77.71 82.64 45.37 42.33 65.64 65.41
ViT-B (scratch) (Dosovitskiy et al., 2021) 75.23 81.86 39.62 45.08 66.18 66.12
ViT-B (Dosovitskiy et al., 2021) 75.47 79.00 50.49 44.83 69.71 66.18
ViT-B (MAE (He et al., 2022) pretrained) 79.97 77.72 56.21 47.07 71.77 61.40
BioCLIP-FT (Stevens et al., 2023b) 79.32 84.06 45.31 42.94 70.07 63.42
@@
Within-Source Splits: The two within-source benchmarks in Table 3 demonstrate higher Macro Recall on the higher-quality split (Train-S2 / Test-S2) for all models except ViT (MAE-pretrained). This trend does not necessarily suggest improved performance due to enhanced data quality, as direct comparisons between Train-S1 and Train-S2 remain inconclusive due to the use of distinct test sets (Test-S1 and Test-S2). Nevertheless, the overall relative performance ranking of models remained consistent across both benchmarks. The stricter quality control applied to Test-S2 samples establishes the (Train-S2 / Test-S2) split as a more reliable benchmark compared to (Train-S1 / Test-S1), as higher-quality test data yields more statistically meaningful measures of performance. In these within-source experiments, we found that BioCLIP-FT notably outperformed other models, performing best on the (Train-S2 / Test-S2) split with 84.06%. This result highlights the potential benefits of leveraging pretrained knowledge from large-scale taxonomic biological data in TreeOfLife-10M (Stevens et al., 2023c).
@@
Loss Function and Class Imbalance. Rare classes exhibit lower recall, an expected outcome given the class imbalance in the training data (see per-class analysis in appendix A.4.3). To mitigate this, we evaluated alternative loss functions. As presented in Table 4, training ViT-L with class-balanced cross-entropy (Cui et al., 2019) did not yield improvements in any split. In contrast, focal loss (Lin et al., 2017) consistently outperformed standard cross-entropy, with further gains observed when using its class-balanced variant (Cui et al., 2019).
@@
CB-CE (Cui et al., 2019) 76.58 76.23 46.96 42.32
Focal (Lin et al., 2017) 78.39 84.33 46.33 41.24
CB-Focal (Cui et al., 2019) 81.86 85.71 52.87 45.65

### v2

@@ 1 Introduction
Coral reefs are among the most biodiverse ecosystems on Earth, providing immense ecological and economic value (barbier2011economic; spalding2017global). Yet these vibrant habitats are increasingly threatened by anthropogenic pressures, including climate change, overfishing, pollution, and ocean acidification (cooley2022coral; bellwood2004coral; hughes2017coral). As these stressors intensify, coral cover and overall reef resilience decline, underscoring the need for effective conservation and restoration (Brandl2019; bellwood2004coral). A critical component of reef conservation is habitat mapping and coverage analysis, which typically relies on expert taxonomists to annotate in situ underwater images (hill2004methods; beijbom2015coralnet). However, this dependence on manual labeling severely limits the scale and speed of monitoring (Brandl2019; bellwood2004coral). Machine learning (ML) offers a path to automation, but progress is impeded by two factors: (i) the scarcity of well-established, ML-ready datasets and benchmarks, and (ii) persistent domain shift, whereby classifiers degrade on sites with different photographic conditions, local assemblages, or label conventions (Belcher2023; chen2021coral).
CoralNet (beijbom2015coralnet; chen2021coral), one of the leading annotation platforms, fine-tunes a global model on user-defined “sources.” However, these source-specific classifiers frequently show poor performance on unseen sites owing to variations in taxonomy, imaging protocols, and local reef composition (williams2019coralnet). ReefCloud.ai (reefcloud) has improved label standardization but remains inaccessible to public researchers. To address these data and accessibility gaps, we introduce ReefNet as a complementary resource, providing a large-scale, publicly available coral reef imagery dataset with point-label annotations taxonomically aligned to the World Register of Marine Species (WoRMS(worms). Figure 1 shows an overview of the dataset’s structure. ReefNet unifies data from 76 carefully selected CoralNet sources and includes an additional $1.3\text{K}$ images that we collected and annotated from the Al-Wajh lagoon in the Red Sea, comprising 4,6094{,}609 expert annotations of Red Sea hard coral genera, offering a valuable contribution from an understudied biogeographic region. In total, the dataset contains $\sim 925\text{K}$ genus-level scleractinian coral annotations (Table 1). Rigorous manual verification by marine biologists was conducted on 8,962 annotations spanning various sources and labels. This process yielded a high-confidence repository of hard coral annotations across diverse biogeographic regions. In addition to point annotations, we enrich ReefNet with textual descriptions for each genus generated from scanned books (coralvol1; coralvol3; staghorn), enabling language-grounded classification and supporting the development of vision-language models (VLMs).
@@ 2 Related Work
Several large-scale machine-learning datasets have recently emerged to facilitate biodiversity classification. One example is TreeOfLife-10M (treeoflife_10m), which consolidates imagery from iNat2021 (iNat2021), BIOSCAN-1M (BIOSCAN), and the Encyclopedia of Life (eol.org), surpassing 10 million images of terrestrial and marine organisms. Despite its scope, underwater taxa remain underrepresented in such general biodiversity repositories, largely because of the substantial logistical challenges associated with acquiring and annotating in situ marine imagery. To address this gap, BenthicNet (BenthicNet) was introduced, aggregating seafloor images from multiple surveys, including the Seaview Survey Photo-quadrat (gonzalez2019seaview) and Reef Life Survey (Edgar2020RLS), among others (SQUIDLE). BenthicNet constitutes a major step toward comprehensive marine coverage, but it primarily supports broad-scale benthic habitat labeling using the CATAMI classification scheme (Althaus2015CATAMI), and it also supports the WORMS (worms)taxonomy as it has 887533 annotations of which 287K are hard coral (’Scleractinia’) annotations.(Table 1). Other coral-specific datasets have been proposed to support tasks such as dense segmentation. For instance, CoralVOS and CoralSCOP (coralvos; coralscop) offer pixel-level masks but generally categorize corals only at the order level (Scleractinia). Consequently, their utility for fine-grained ecological studies is limited. In contrast, MosaicsUCSD and Eilat (edwards2017; Beijbom2016; Alonso2019) provide annotations at genus or species levels but are geographically constrained to a single reef site, reducing their ability to generalize across regions. Meanwhile, CoralNet (beijbom2015coralnet) hosts a large volume of benthic imagery with both human- and machine-generated labels, but does not typically map them to an accurate, globally recognized taxonomy such as the WoRMS. Coralscapes (sauder2025coralscapesdatasetsemanticscene) extends benthic classification to a sizeable Red Sea dataset with over 170K polygon annotations, focusing on habitat cover types rather than genus- or species-level coral identification. ReefNet builds on prior efforts while addressing several key limitations, providing one of the most comprehensive collections of expert-verified coral annotations available. These annotations are mapped to the genus level across a diverse, global set of reef sites, addressing challenges like limited taxonomic resolution, geographic bias, and inconsistent labeling standards. By aligning its taxonomy with the World Register of Marine Species (WoRMS), ReefNet ensures consistency and compatibility with ongoing coral biodiversity research, supporting a broad spectrum of ecological and machine learning tasks from habitat classification to genus-level analysis. We provide further clarification of our advantages over CoralNet in appendix A.1.
@@
Table 1: Comparison With Existing Coral Classification Datasets. *Image counts reflect availability as of May 13, 2025. Hard coral/genus-level annotations are not specified for CoralNet, since its labels are not standardized. MosaicsUCSD includes 16 annotated orthomosaics, each from $\sim$1.5K images. CoralSCOP (coralscop) contains 330,144 binary coral/non-coral masks (not limited to hard corals). CoralVOS (coralvos) likewise provides only binary coral masks, so a hard coral count is N/A.
@@
CoralNet (beijbom2015coralnet) World Sparse points Species No 4,524,792* N/A No
CoralSCOP (coralscop) World Masks Order No 41,297 330,144 No
CoralVOS (coralvos) 17 sites (South China Sea) Masks Order No 60,456 N/A No
MosaicsUCSD (edwards2017) Palmyra Masks Species No 16 44,008 Yes
Eilat (Raphael2020eilat) Eilat Sparse points Genus No 212 $\sim$12,000 Yes
BenthicNet (BenthicNet) World
@@
Coralscapes (sauder2025coralscapesdatasetsemanticscene) Red Sea (5 countries) Masks N/A No 2,075 N/A No
@@ 3.2 Label Standardization and Mapping
To ensure taxonomic traceability and biological consistency, we mapped annotation labels to the World Register of Marine Species (WoRMS worms). This step was applied, where applicable, to annotations from the 76 curated CoralNet sources and the Al-Wajh lagoon dataset. Hard coral labels were manually aligned with canonical scientific names and corresponding AphiaIDs [^1] An AphiaID is a unique identifier from the WoRMS database, linking each label to its corresponding entry in the hierarchical taxonomy. As coral taxonomy evolves, AphiaIDs allow users to verify the current validity of labels.. This mapping enabled consistent aggregation of biological entities across sources and ensured compatibility with other biodiversity databases. The initial mapping revealed $\sim$920k annotations at multiple taxonomic levels, consolidated to genus-level hard corals. The dataset covers 44 reef-building hard coral (Order: Scleractinia) genera spanning 20 families. However, this initial label space exhibited inconsistencies in naming, taxonomic granularity, and semantic intent, necessitating further filtering and restructuring.
@@
Label Filtering and Semantic Consolidation. To construct a benchmark dataset suitable for training generalizable AI models, we applied a biologically informed filtering strategy targeting syntactic variation (e.g., “Staghorn coral” vs. Acropora cervicornis) and taxonomic inconsistency (e.g., mixing species-, genus-, and family-level labels). This step was motivated by the inherent difficulty of fine-grained coral identification from imagery alone (chen2021coral; BenthicNet), which often results in community datasets with inconsistent or imprecise taxonomic labeling.
@@
Figure 3: Geographic Distribution of the Annotations. Sources located within two degrees of Latitude and Longitude are grouped into a single point. One source did not contain any location data and is displayed in Antarctica. Marine Ecoregions of the World are shown in light blue (spalding_ecoregions).
@@ 3.3.1 Annotation Distribution
The ReefNet dataset comprises 334,162 images, of which 181,223 contain hard coral annotations, totaling 924,626 point annotations. These annotations span 26 marine ecoregions across the tropical belt sensu (spalding_ecoregions) (see appendix A.2.5 for details). The distribution is uneven: the Hawaiian Ecoregion alone contributes over 221K annotations, followed by the Samoa Islands (202K) and the Mariana Islands (108K). In contrast, regions such as the Southwestern Caribbean, Floridian, and Eastern Brazil are sparsely represented, each with fewer than 8K samples. This geographic imbalance reflects broader disparities in coral monitoring efforts and may influence model generalization to underrepresented areas. Among all annotations, 921,351 are assigned to one of 44 classes, capturing a wide range of coral diversity. The most frequent genera, Porites, Montipora, Acropora, and Pocillopora, dominate shallow reefs globally (Figure 2, Figure 4), while several genera appear in less than 1K annotations, illustrating the dataset’s imbalanced distribution.
@@ 4 ReefNet Benchmarks
A key practical challenge for coral-reef AI is the domain gap encountered when models trained on imagery from one dataset (a “source”) are applied to imagery collected by another dataset. Models trained on datasets from specific locations frequently exhibit poor generalization to ecologically or geographically distinct areas (williams2019coralnet; Belcher2023; wyatt2025safeai). Variations in camera setups, water clarity, depth, and annotation protocols, even for the same taxa, can degrade performance. To systematically evaluate model robustness and practical applicability in genus-level coral classification, we define two primary benchmarking scenarios: (i) Within-source fine-tuning: Images from a single source (Source A) are partitioned into distinct training, validation, and test subsets. Models are trained on a labeled subset and evaluated on the remainder. Although used by platforms such as CoralNet and ReefCloud, this method requires site-specific manual annotations, limiting scalability. (ii) Cross-source deployment: Models trained on all available images from Source A are directly evaluated on a different source (Source B), without further fine-tuning. This setting assesses generalization across domains, highlighting the model’s ability to overcome domain shifts. ReefNet provides benchmarks for both settings with different training and testing configurations. A quantitative overview of all the splits is in Table 2.
@@ 4.1 ReefNet Within‑Source Split
For the within-source setting, we consider two splits that share an identical label set but with different expert agreement quality. In each split, each source is divided on the image level to prevent information leakage using Multilabel Stratified Shuffle Split (sechidis2011stratification; brady2017iterative). The two data‑quality variants considered: (i) Train-S1 / Test-S1, with no source–genus filtration, of 81 % expert agreement, 802,956 training annotations from 70 sources; 9,972 validation annotations from 64 sources; 40,881 test annotations from 69 sources. (ii) Train-S2 / Test-S2, with source–genus filtration, of 92 % expert agreement percentage, 445,985 training annotations from 68 sources; 9,999 validation annotations from 63 sources; 40,881 test annotations from 66 sources.
@@ 5.1 Experimental setup
Evaluated Models. We evaluate three types of models across all benchmarks: (i) Large Scale Pre-trained models fine-tuned on our dataset (tan2019efficientnet; chen2021coral; he2016deep; liu2022convnet; vit; bao2021beit; swin; deit), (ii) Vision-Language Models (VLMs) evaluated in a zero-shot setting (openclip; clip; BioCLIP; siglip), and (iii) Multimodal Large Language Models (MLLMs) also evaluated in a zero-shot setting (qwen25). Model performance is evaluated using Macro Recall, sometimes referred to as the balanced accuracy score (scikit-learn). Macro Recall is particularly suited for our task as the ReefNet data shows a high degree of class imbalance 2. Additionally, for top-performing models, per-class recall, precision, and F1 scores are provided in appendices( A.4.3A.4.5). Additionally, we fine-tune BioCLIP on our hierarchical dataset, down to the Genus level (referred to as BioCLIP-FT), to investigate its performance on hard coral hierarchical classification, as it reported good performance on other datasets spanning plants, animals, and fungi on TreeOfLife-10M (treeoflife_10m). The details about the training setup for all models have been included in the supplementary material.
@@ 5.2 Finetuning Results
EfficientNet (tan2019efficientnet) 78.36 82.41 48.01 41.78 69.29 64.59
ResNet (he2016deep) 63.91 64.49 37.96 28.49 64.14 51.22
ConvNext (liu2022convnet) 77.52 82.88 42.21 47.06 69.19 68.08
BEiT (bao2021beit) 72.18 73.63 45.56 32.17 64.24 62.31
DeiT (deit) 74.36 77.40 45.46 40.94 63.19 66.67
Swin (swin) 77.71 82.64 45.37 42.33 65.64 65.41
ViT-B (scratch) (vit) 75.23 81.86 39.62 45.08 66.18 66.12
ViT-B (vit) 75.47 79.00 50.49 44.83 69.71 66.18
ViT-B (MAE (he2022masked) pretrained) 79.97 77.72 56.21 47.07 71.77 61.40
BioCLIP-FT (bioclip2023) 79.32 84.06 45.31 42.94 70.07 63.42
@@
Within-Source Splits: The two within-source benchmarks in Table 3 demonstrate higher Macro Recall on the higher-quality split (Train-S2 / Test-S2) for all models except ViT (MAE-pretrained). This trend does not necessarily suggest improved performance due to enhanced data quality, as direct comparisons between Train-S1 and Train-S2 remain inconclusive due to the use of distinct test sets (Test-S1 and Test-S2). Nevertheless, the overall relative performance ranking of models remained consistent across both benchmarks. The stricter quality control applied to Test-S2 samples establishes the (Train-S2 / Test-S2) split as a more reliable benchmark compared to (Train-S1 / Test-S1), as higher-quality test data yields more statistically meaningful measures of performance. In these within-source experiments, we found that BioCLIP-FT notably outperformed other models, performing best on the (Train-S2 / Test-S2) split with 84.06%. This result highlights the potential benefits of leveraging pretrained knowledge from large-scale taxonomic biological data in TreeOfLife-10M (treeoflife_10m).
@@
Loss Function and Class Imbalance. Rare classes exhibit lower recall, an expected outcome given the class imbalance in the training data (see per-class analysis in appendix A.4.3). To mitigate this, we evaluated alternative loss functions. As presented in Table 4, training ViT-L with class-balanced cross-entropy (Cui2019ClassBalancedLB) did not yield improvements in any split. In contrast, focal loss (focal_loss) consistently outperformed standard cross-entropy, with further gains observed when using its class-balanced variant (Cui2019ClassBalancedLB).
@@
CB-CE (Cui2019ClassBalancedLB) 76.58 76.23 46.96 42.32
Focal (focal_loss) 78.39 84.33 46.33 41.24
CB-Focal (Cui2019ClassBalancedLB) 81.86 85.71 52.87 45.65